# Cross-View Geo-Localization with Street-View and VHR Satellite Imagery in Decentrality Settings

### v1

@@ 3.3 Comparison with Previous Datasets
CVUSA CVACT VIGOR CVGlobal CVSat
@@
Table 3: Comparison between the proposed CVSat dataset and existing datasets for cross-view geo-localization task.
Figure 3: Visualization of decentrality conditions across different datasets. The red star in CVUSA, CVACT and CVGlobal represents center alignment, while the yellow box in VIGOR and the red box in CVSat represent the hit areas, indicating different degrees of decentrality.
Table 3 provides a detailed comparison between our dataset and previous datasets CVUSA [10], CVACT [11], VIGOR [12] and CVGlobal [21]. We also visualize the decentrality conditions in existing datasets alongside our proposed CVSat dataset for comparison, as shown in Figure 3. In the CVUSA [10], CVACT [11] and CVGlobal [21] datasets, decentrality is overlooked due to the simple center-aligned settings, which are inadequate for real-world applications where perfectly matched reference images cannot be guaranteed. In the VIGOR [12] dataset, due to a high overlap of 50% and dataset balancing, the best-matched image pairs are mostly confined to the yellow box, indicating a limited portion of decentrality.
The statistical analysis results of VIGOR are also presented in Table 4. While VIGOR attempts to define a hit area for positive samples, a small proportion of the best-matched image pairs (positive samples) are erroneously assigned to subsets with high decentrality. Additionally, although the dataset includes ”semi-positive” street-view images with larger decentrality, there is always an alternative reference image available to establish a ”positive” pair for each street-view image. Consequently, the VIGOR dataset is inadequate for effectively investigating the decentrality issue, even with a reorganization of its original structure.
@@
Compared to previous datasets, our proposed CVSat dataset can better cover the issue of decentrality, providing a more comprehensive testbed for CVGL. In addition to introducing the issue of decentrality, CVSat covers a larger scope and a wider variety of landscapes. As illustrated in Figure 2(b) and Figure 4, CVSat encompasses urban and suburban areas. For example, the coverage of Chicago in CVSat (619 $km^{2}$) is almost 20 times larger than in VIGOR (30.3 $km^{2}$). CVSat includes samples from eight cities worldwide, each exceeding 400 $km^{2}$. This makes CVSat the largest and most realistic dataset for cross-view geo-localization.
Figure 4: Aerial images of eight cities with diverse landscapes from across the world and the distributions of panoramas (red dots) in the CVSat dataset.
@@ 3.4 Data Collection
As shown in Figure 4, the CVSat dataset covers over 400 $km^{2}$ in each of eight diverse cities around the world. The dataset consists of 422,760 aerial images sourced from Esri World Imagery [41], captured at zoom level 18 with a ground resolution of approximately 0.597 $m$. Each aerial image has a resolution of $224\times 224$ pixels. Additionally, CVSat features 174,934 street-view panoramas obtained using the Google Street View. These panoramas are randomly distributed within the coverage area of the aerial images, with an average interval of about 500 $m$ between samples. The panoramas are North-aligned, and each has a resolution of $2048\times 1024$ pixels. To comply with Google Street View’s policies, we will provide only the panorama IDs and instructions for downloading them, enabling users to access the dataset independently.
@@ 3.5 The Evaluation Protocol
We adopt the SAME and CROSS settings as used in VIGOR. In the SAME, the model is trained with training data from all cities and tested with testing data from all cities. In the CROSS setting, the model is trained with data from four cities (Chicago, London, Sydney, Tokyo) and tested with data from the remaining four cities (Johannesburg, Rio, Taipei, Vancouver). Furthermore, a novel evaluation is proposed within the testing data of CVSat to assess the model’s robustness for cross-view geo-localization with increasing decentrality.
@@ 4.2 AuxGeo Model
Given a cross-view image pair $(I^{s},I^{a})$, the fine-grained features generated by the backbone are denoted as $(F^{s},F^{a})$, where $F^{s}\in\mathbb{R}^{H^{s}\times W^{s}\times C}$ and $F^{a}\in\mathbb{R}^{H^{a}\times W^{a}\times C}$. Here, $H$, $W$, and $C$ represent height, width, and channels, respectively. Existing methods achieve superior representations $(f^{s},f^{a})$, where $f^{s},f^{a}\in\mathbb{R}^{C}$, from the fine-grained features using self-attention [38, 42] or feature recombination [15]. However, it has been demonstrated that these capabilities can be replaced by a robust backbone with a simple global average pooling [14]. Therefore, moving beyond current methods, our proposed AuxGeo focuses on leveraging information that cannot be directly included in the representation but can enhance the backbone’s capabilities.
@@ 4.2.1 Model Overview
The proposed AuxGeo incorporates a multi-metric optimization network, which includes a contrastive representation learning structure derived from Sample4Geo [14] and two auxiliary modules, as depicted in Figure 5(A1).
@@ 4.2.2 BEV Intermediary Module
The BIM (BEV Intermediary Module) uses the BEV-view image $I^{bev}$ generated from the street-view panorama $I^{s}$ through a geometry-only transformation introduced by Wang et al. [43] as its input. While the BEV image effectively overcomes the cross-view problem under decentrality conditions, it only retains ground-level details (like road markings) and omits others (such as houses and trees) because it’s a subset of the street-view panorama. Therefore, we incorporate BEV-view images as intermediaries in our method rather than using them directly as query images.
@@ 4.2.3 Position Constraint Module
The Position Constraint Module (PCM) processes the coarse-grained features $f^{s}$ and $f^{bev}$ alongside the fine-grained feature $F^{a}$, using position prior information as constraints to enhance the backbone’s performance involving decentrality, as depicted in Figure 5(A2). The position prior $Pos$, indicating the location of the query image on the reference image, can be easily determined using geo-tags. Unlike previous offset regression methods that apply the position prior only at the coarse-grained level [12], we recognize that the aerial-view image, acting as the map, typically has a larger coverage compared to the street-view image as the query. Therefore, we designed a multi-grained position regression task within the PCM.
@@
The coarse-grained features $f^{s}$ and $f^{bev}$ are utilized to compute the similarity with each pixel of the fine-grained feature $F^{a}$, producing similarity maps $M^{s}$ and $M^{bev}$. This process does not involve any additional learnable components to maintain the backbone’s representation purity. From these similarity maps, the regressed positions $Reg^{s}$ and $Reg^{bev}$ are derived through a soft matching process, as depicted in Figure 7. Considering the variation in coverage ratio between cross-view images in different scenarios (for example, the aerial-view image in CVSat has a larger coverage than that in VIGOR, while the street-view image always has a similar coverage), a feature pyramid is employed in the PCM to utilize the position prior at different scales. The original fine-grained feature is processed at multiple scales: $lv1$ for the original feature, $lv2$ for 2x down-sampled features, and $lv3$ for 4x down-sampled features, each generated using average pooling. In this setup, position priors at different scales are denoted as $Pos_{lvi}$, and the regressed positions at different levels are denoted as $Reg^{s}_{lvi}$ and $Reg^{bev}_{lvi}$ , obtained through a soft matching process from the similarity maps. To supervise the network during training, an L2 Euclidean distance loss, denoted as $d(\cdot,\cdot)$, is employed. $N$ represents the number of layers in the feature pyramid, and $N=3$. The loss function for the PCM is expressed as follows:
@@ 5.1.1 Datasets
We evaluate AuxGeo on four cross-view geo-localization datasets: CVUSA, CVACT, VIGOR, and our proposed CVSat dataset. CVUSA and CVACT employ a center-aligned, one-to-one retrieval paradigm between cross-view images, whereas VIGOR and CVSat extend beyond one-to-one retrieval, featuring non-center-aligned settings. Notably, CVSat supports cross-view geo-localization in a significantly larger and seamless environment, encompassing diverse styles of cities worldwide.
  • 1.
@@
  • 4.
    CVSat includes 174,934 street images and 422,760 aerial images from eight cities worldwide. The reference images provide a much larger and seamless coverage, without center-aligned settings. The dataset spans urban, suburban, and rural areas.
@@ 5.1.2 Evaluation Metrics
Following previous works, we use the R@$K,K=\left\{{1,5,10,1\%}\right\}$ as the evaluation metric, which represents the ratio of correct retrievals within the top $K$ results. Additionally, for VIGOR and CVSat, we report the hit rate, indicating the probability that the top-1 retrieved reference image covers the query image.
@@ 5.1.3 Implementation Details
We employ a ConvNeXt-B backbone [44] pretrained on ImageNet-22K. Following previous works, we use an input resolution of $384\times 768$ for street-view images and $384\times 384$ for aerial-view images. In Equation (4), the hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 0.1, 0.1, and 0.05, respectively. Models are trained on a server equipped with 8 NVIDIA V100-32G GPUs.
@@ 5.2 Comparison with State-of-the-art Methods
We comapre our AuxGeo with several state-of-the-art methods, including TransGeo [38], GeoDTR [26], FRGeo [15] and our baseline Sample4Geo [14].
Method Backbone CVSat VIGOR
@@
Table 5: Quantitative comparison between our AuxGeo and current state-of-the-art methods on VIGOR and CVSat. * indicates that the backbone used for FRGeo on CVSat is ConvNeXt-B for fair comparison, while on VIGOR, ConvNeXt-T is used as reported in the original paper. The best results are shown in bold.
5.2.1 Results on VIGOR and CVSat
Beyond the center-aligned setting, we report the cross-view geo-localization performance on VIGOR and CVSat, with results shown in Table 5. On VIGOR, AuxGeo achieves state-of-the-art performance. This improvement is attributed to the enhanced capability of the backbone network, activated by the BIM and PCM modules in AuxGeo. On CVSat, performance significantly decreases due to the greater decentrality. Despite this, AuxGeo can also outperform previous state-of-the-art methods, improving the R@1 by 3.30% in SAME and 4.21% in CROSS compared to the previous state-of-the-art method, Sample4Geo.
@@ 5.2.1 Results on VIGOR and CVSat
Table 6: Quantitative comparison of R@1 score on subsets of CVSat SAME, considering increasing degrees of decentrality. The best results are shown in bold.
@@ 5.2.2 Results on Increasing Decentrality
To further validate the effectiveness of mitigating the decentrality problem, we conducted a series of experiments. Models were trained using data with various degrees of decentrality and evaluated across Subsets 1 to 4 within the testing data. The results, presented in Table 6, reveal the following: a) Geo-localization performance significantly degrades with increasing decentrality. The performance of both the Sample4Geo and AuxGeo on Subsets 1 and 2 is comparable to that on the VIGOR dataset, where decentrality is similar to that of Subsets 1 and 2 in CVSat. However, performance on Subsets 3 and 4 experiences a sharp decline for both methods, indicating that decentrality poses a substantial challenge, warranting further investigation. b) AuxGeo, despite experiencing a performance decrease, demonstrates progressively better improvement compared to the baseline. This suggests that the proposed modules effectively contribute to overcoming the challenges induced by decentrality.
@@
SAME VIGOR $\longrightarrow$ CVSat
@@
SAME CVSat $\longrightarrow$ VIGOR
@@
Table 7: Generalization performance when trained on VIGOR and evaluated on CVSat and vice versa. The best results are shown in bold.
@@ 5.2.3 Results on Generalization Capabilities
Besides the evaluation of the CROSS settings in VIGOR and CVSat, we also evaluated the generalization capabilities of models trained on VIGOR and tested on CVSat, and vice versa, to test the transferability between images from different sources. The results, shown in Table 7, demonstrate that AuxGeo achieves better performance. This advantage is particularly emphasized in the generalization capabilities from larger datasets (CVSat) to smaller ones (VIGOR), which indicates the importance of the auxillary contents provided by modules in AuxGeo.
@@ 6.1 Effectiveness of Components
To evaluate the effectiveness of the proposed modules (BIM and PCM) in the AuxGeo method, we conducted a series of ablation experiments. For fair comparison, the settings of strategies during training and testing remained consistent across all configurations. The results on VIGOR and CVSat datasets are presented in Table 9. The results demonstrate that integrating either BIM or PCM yields improvements, with the best performance achieved when both BIM and PCM are incorporated. These findings offer a new perspective on cross-view geo-localization. As an image retrieval-based method, it demonstrates that certain elements, which can’t be directly used during inference (such as partial viewpoint transform and fine-grained cross-view feature correspondences), can be effectively utilized as auxiliary components during training. This approach maximizes the use of these elements to better train the backbone model.
@@
Figure 8: Heatmap visualization of CVSat dataset images generated by baseline (Sample4Geo) and AuxGeo models. The third column indicates the ground truth position (red dot) of the street-view image on the corresponding aerial-view image.
Figure 9: Visualization results of the cross-view geo-localization in CVSat dataset.

### v2

@@ 3.3 Comparison with Previous Datasets
CVUSA CVACT VIGOR CVGlobal DReSS
@@
Table 3: Comparison between the proposed DReSS dataset and existing datasets for cross-view geo-localization task.
Figure 3: Visualization of decentrality conditions across different datasets. The red star in CVUSA, CVACT and CVGlobal represents center alignment, while the yellow box in VIGOR and the red box in DReSS represent the hit areas, indicating different degrees of decentrality.
Table 3 provides a detailed comparison between our dataset and previous datasets CVUSA (Zhai et al., 2017), CVACT (Liu and Li, 2019), VIGOR (Zhu et al., 2021) and CVGlobal (Ye et al., 2025). We also visualize the decentrality conditions in existing datasets alongside our proposed DReSS dataset for comparison, as shown in Figure 3. In the CVUSA (Zhai et al., 2017), CVACT (Liu and Li, 2019) and CVGlobal (Ye et al., 2025) datasets, decentrality is overlooked due to the simple center-aligned settings, which are inadequate for real-world applications where perfectly matched reference images cannot be guaranteed. In the VIGOR (Zhu et al., 2021) dataset, due to a high overlap of 50% and dataset balancing, the best-matched image pairs are mostly confined to the yellow box, indicating a limited portion of decentrality.
The statistical analysis results of VIGOR are also presented in Table 4. While VIGOR attempts to define a hit area for positive samples, a small proportion of the best-matched image pairs (positive samples) are erroneously assigned to subsets with high decentrality. Additionally, although the dataset includes "semi-positive" street-view images with larger decentrality, there is always an alternative reference image available to establish a "positive" pair for each street-view image. Consequently, the VIGOR dataset is inadequate for effectively investigating the decentrality issue, even with a reorganization of its original structure.
@@
Compared to previous datasets, our proposed DReSS dataset can better cover the issue of decentrality, providing a more comprehensive testbed for CVGL. In addition to introducing the issue of decentrality, DReSS covers a larger scope and a wider variety of landscapes. As illustrated in Figure 2(b) and Figure 4, DReSS encompasses urban and suburban areas. For example, the coverage of Chicago in DReSS (619 $km^{2}$) is almost 20 times larger than in VIGOR (30.3 $km^{2}$). DReSS includes samples from eight cities worldwide, each exceeding 400 $km^{2}$. This makes DReSS the largest and most realistic dataset for cross-view geo-localization.
Figure 4: Aerial images of eight cities with diverse landscapes from across the world and the distributions of panoramas (red dots) in the DReSS dataset.
@@ 3.4 Data Collection
As shown in Figure 4, the DReSS dataset covers over 400 $km^{2}$ in each of eight diverse cities around the world. The dataset consists of 422,760 aerial images sourced from Esri World Imagery (Esri, 2024), captured at zoom level 18 with a ground resolution of approximately 0.597 $m$. Each aerial image has a resolution of $224\times 224$ pixels. Additionally, DReSS features 174,934 street-view panoramas obtained using the Google Street View. These panoramas are randomly distributed within the coverage area of the aerial images, with an average interval of about 500 $m$ between samples. The panoramas are North-aligned, and each has a resolution of $2048\times 1024$ pixels. To comply with Google Street View’s policies, we will provide only the panorama IDs and instructions for downloading them, enabling users to access the dataset independently.
@@ 3.5 The Evaluation Protocol
We adopt the SAME and CROSS settings as used in VIGOR. In the SAME, the model is trained with training data from all cities and tested with testing data from all cities. In the CROSS setting, the model is trained with data from four cities (Chicago, London, Sydney, Tokyo) and tested with data from the remaining four cities (Johannesburg, Rio, Taipei, Vancouver). Furthermore, a novel evaluation is proposed within the testing data of DReSS to assess the model’s robustness for cross-view geo-localization with increasing decentrality.
@@ 4.2 AuxGeo Model
Given a cross-view image pair $(I^{s},I^{a})$, the fine-grained features generated by the backbone are denoted as $(F^{s},F^{a})$, where $F^{s}\in\mathbb{R}^{H^{s}\times W^{s}\times C}$ and $F^{a}\in\mathbb{R}^{H^{a}\times W^{a}\times C}$. Here, $H$, $W$, and $C$ represent height, width, and channels, respectively. Existing methods achieve superior representations $(f^{s},f^{a})$, where $f^{s},f^{a}\in\mathbb{R}^{C}$, from the fine-grained features using self-attention (Zhu et al., 2022, 2023) or feature recombination (Zhang and Zhu, 2024). However, it has been demonstrated that these capabilities can be replaced by a robust backbone with a simple global average pooling (Deuser et al., 2023). Therefore, moving beyond current methods, our proposed AuxGeo focuses on leveraging information that cannot be directly included in the representation but can enhance the backbone’s capabilities.
@@ 4.2.1 Model Overview
The proposed AuxGeo incorporates a multi-metric optimization network, which includes a contrastive representation learning structure derived from Sample4Geo (Deuser et al., 2023) and two auxiliary modules, as depicted in Figure 5(A1).
@@ 4.2.2 BEV Intermediary Module
The BIM (BEV Intermediary Module) uses the BEV-view image $I^{bev}$ generated from the street-view panorama $I^{s}$ through a geometry-only transformation introduced by Wang et al. (Wang et al., 2024) as its input. While the BEV image effectively overcomes the cross-view problem under decentrality conditions, it only retains ground-level details (like road markings) and omits others (such as houses and trees) because it’s a subset of the street-view panorama. Therefore, we incorporate BEV-view images as intermediaries in our method rather than using them directly as query images.
@@ 4.2.3 Position Constraint Module
The Position Constraint Module (PCM) processes the coarse-grained features $f^{s}$ and $f^{bev}$ alongside the fine-grained feature $F^{a}$, using position prior information as constraints to enhance the backbone’s performance involving decentrality, as depicted in Figure 5(A2). The position prior $Pos$, indicating the location of the query image on the reference image, can be easily determined using geo-tags. Unlike previous offset regression methods that apply the position prior only at the coarse-grained level (Zhu et al., 2021), we recognize that the aerial-view image, acting as the map, typically has a larger coverage compared to the street-view image as the query. Therefore, we designed a multi-grained position regression task within the PCM.
@@
The coarse-grained features $f^{s}$ and $f^{bev}$ are utilized to compute the similarity with each pixel of the fine-grained feature $F^{a}$, producing similarity maps $M^{s}$ and $M^{bev}$. This process does not involve any additional learnable components to maintain the backbone’s representation purity. From these similarity maps, the regressed positions $Reg^{s}$ and $Reg^{bev}$ are derived through a soft matching process, as depicted in Figure 7. Considering the variation in coverage ratio between cross-view images in different scenarios (for example, the aerial-view image in DReSS has a larger coverage than that in VIGOR, while the street-view image always has a similar coverage), a feature pyramid is employed in the PCM to utilize the position prior at different scales. The original fine-grained feature is processed at multiple scales: $lv1$ for the original feature, $lv2$ for 2x down-sampled features, and $lv3$ for 4x down-sampled features, each generated using average pooling. In this setup, position priors at different scales are denoted as $Pos_{lvi}$, and the regressed positions at different levels are denoted as $Reg^{s}_{lvi}$ and $Reg^{bev}_{lvi}$ , obtained through a soft matching process from the similarity maps. To supervise the network during training, an L2 Euclidean distance loss, denoted as $d(\cdot,\cdot)$, is employed. $N$ represents the number of layers in the feature pyramid, and $N=3$. The loss function for the PCM is expressed as follows:
@@ 5.1.1 Datasets
We evaluate AuxGeo on four cross-view geo-localization datasets: CVUSA, CVACT, VIGOR, and our proposed DReSS dataset. CVUSA and CVACT employ a center-aligned, one-to-one retrieval paradigm between cross-view images, whereas VIGOR and DReSS extend beyond one-to-one retrieval, featuring non-center-aligned settings. Notably, DReSS supports cross-view geo-localization in a significantly larger and seamless environment, encompassing diverse styles of cities worldwide.
@@
  • DReSS includes 174,934 street images and 422,760 aerial images from eight cities worldwide. The reference images provide a much larger and seamless coverage, without center-aligned settings. The dataset spans urban, suburban, and rural areas.
@@ 5.1.2 Evaluation Metrics
Following previous works, we use the R@$K,K=\left\{{1,5,10,1\%}\right\}$ as the evaluation metric, which represents the ratio of correct retrievals within the top $K$ results. Additionally, for VIGOR and DReSS, we report the hit rate, indicating the probability that the top-1 retrieved reference image covers the query image.
@@ 5.1.3 Implementation Details
We employ a ConvNeXt-B backbone (Liu et al., 2022) pretrained on ImageNet-22K. Following previous works, we use an input resolution of $384\times 768$ for street-view images and $384\times 384$ for aerial-view images. In Equation (4), the hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 0.1, 0.1, and 0.05, respectively. Models are trained on a server equipped with 8 NVIDIA V100-32G GPUs, using a batchsize of 128.
@@ 5.2 Comparison with State-of-the-art Methods
We comapre our AuxGeo with several state-of-the-art methods, including TransGeo (Zhu et al., 2022), GeoDTR (Zhang et al., 2023), FRGeo (Zhang and Zhu, 2024) and our baseline Sample4Geo (Deuser et al., 2023).
Method Backbone DReSS VIGOR
@@
Table 5: Quantitative comparison between our AuxGeo and current state-of-the-art methods on VIGOR and DReSS. * indicates that the backbone used for FRGeo on DReSS is ConvNeXt-B for fair comparison, while on VIGOR, ConvNeXt-T is used as reported in the original paper. The best results are shown in bold.
5.2.1 Results on VIGOR and DReSS
Beyond the center-aligned setting, we report the cross-view geo-localization performance on VIGOR and DReSS, with results shown in Table 5. On VIGOR, AuxGeo achieves state-of-the-art performance. This improvement is attributed to the enhanced capability of the backbone network, activated by the BIM and PCM modules in AuxGeo. On DReSS, performance significantly decreases due to the greater decentrality. Despite this, AuxGeo can also outperform previous state-of-the-art methods, improving the R@1 by 3.30% in SAME and 4.21% in CROSS compared to the previous state-of-the-art method, Sample4Geo.
@@ 5.2.1 Results on VIGOR and DReSS
Table 6: Quantitative comparison of R@1 score on subsets of DReSS SAME, considering increasing degrees of decentrality. The best results are shown in bold.
@@ 5.2.2 Results on Increasing Decentrality
To further validate the effectiveness of mitigating the decentrality problem, we conducted a series of experiments. Models were trained using data with various degrees of decentrality and evaluated across Subsets 1 to 4 within the testing data. The results, presented in Table 6, reveal the following: a) Geo-localization performance significantly degrades with increasing decentrality. The performance of both the Sample4Geo and AuxGeo on Subsets 1 and 2 is comparable to that on the VIGOR dataset, where decentrality is similar to that of Subsets 1 and 2 in DReSS. However, performance on Subsets 3 and 4 experiences a sharp decline for both methods, indicating that decentrality poses a substantial challenge, warranting further investigation. b) AuxGeo, despite experiencing a performance decrease, demonstrates progressively better improvement compared to the baseline. This suggests that the proposed modules effectively contribute to overcoming the challenges induced by decentrality.
@@
SAME VIGOR $\longrightarrow$ DReSS
@@
SAME DReSS $\longrightarrow$ VIGOR
@@
Table 7: Generalization performance when trained on VIGOR and evaluated on DReSS and vice versa. The best results are shown in bold.
@@ 5.2.3 Results on Generalization Capabilities
Besides the evaluation of the CROSS settings in VIGOR and DReSS, we also evaluated the generalization capabilities of models trained on VIGOR and tested on DReSS, and vice versa, to test the transferability between images from different sources. The results, shown in Table 7, demonstrate that AuxGeo achieves better performance. This advantage is particularly emphasized in the generalization capabilities from larger datasets (DReSS) to smaller ones (VIGOR), which indicates the importance of the auxillary contents provided by modules in AuxGeo.
@@ 6.1 Effectiveness of Components
To evaluate the effectiveness of the proposed modules (BIM and PCM) in the AuxGeo method, we conducted a series of ablation experiments. For fair comparison, the settings of strategies during training and testing remained consistent across all configurations. The results on VIGOR and DReSS datasets are presented in Table 9. The results demonstrate that integrating either BIM or PCM yields improvements, with the best performance achieved when both BIM and PCM are incorporated. These findings offer a new perspective on cross-view geo-localization. As an image retrieval-based method, it demonstrates that certain elements, which can’t be directly used during inference (such as partial viewpoint transform and fine-grained cross-view feature correspondences), can be effectively utilized as auxiliary components during training. This approach maximizes the use of these elements to better train the backbone model.
@@
Figure 8: Heatmap visualization of DReSS dataset images generated by baseline (Sample4Geo) and AuxGeo models. The third column indicates the ground truth position (red dot) of the street-view image on the corresponding aerial-view image.
Figure 9: Visualization results of the cross-view geo-localization in DReSS dataset.